OpenStudy (anonymous):

An elephant’s tusks grow throughout its life. Each month, an elephant tusk grows about 1 cm. Suppose you started observing an elephant when its tusks were 12 cm long. Write an equation in point-slope form that describes the length of the elephant’s tusks after m months of observation.

OpenStudy (anonymous):

I need help!

OpenStudy (anonymous):

y-y = m(x-x) point slope

OpenStudy (anonymous):

How do you solve it using that formula?

OpenStudy (anonymous):

Give me a sec

OpenStudy (anonymous):

Okay.

OpenStudy (danjs):

They give you the rate of change, the slope m as 1 cm/month They also give you one data point, when time is 0 months , and length is 12 cm (x,y)=(0,12)

OpenStudy (anonymous):

So, what will the equation be?

OpenStudy (anonymous):

I still don't understand.

OpenStudy (anonymous):

@Brian42 do you know?

OpenStudy (anonymous):

y-12 = x

OpenStudy (anonymous):

or Y-12=1(x-0)

OpenStudy (danjs):

slope m=1 point (x1,y1)=(0,12) y-y1=m*(x-x1)

OpenStudy (anonymous):

thats the formula for the point slope form of a line I believe the equation to find the length of the elephant's tusk is y-12=1(x-0), which is the same as y-12=x, where is the number of months

OpenStudy (anonymous):

x*

OpenStudy (anonymous):

Thanks, I have another question: What is the point-slope form of a line with slope 3/ 4x that passes through the point (16, 5)?

OpenStudy (anonymous):

after 1 month y-12 = 1 which means the elephant's tusk would be 13 cm after 2 months y-12 = 2 which would be 14 cm

OpenStudy (anonymous):

(Y-5) = 3/4(x-16)

OpenStudy (anonymous):

Show me your work.

OpenStudy (anonymous):

lol (16,5) = (X,Y) and the point slope form is Y-Y1 = M(X-X1) M = the slope 3/4

OpenStudy (anonymous):

One last question, Use the point-slope form of the equation y 6 8(x 1). What is the y-value of the y-intercept?

OpenStudy (anonymous):

you have to retype that. Not sure what that is.

OpenStudy (anonymous):

*y-6=8(x+1)

OpenStudy (anonymous):

to find the y intercept set x = 0 Y-6 = 8x+8 Y = 8x+14

OpenStudy (anonymous):

Thanks so so much!!